Find an equation of a circle having a diameter with endpoints (-3,5) and (7,3)
Answer by stanbon(75887)

Find an equation of a circle having a diameter with endpoints (-3,5) and (7,3)
1st: Find the center: (x,y)
x = (-3+7)/2 = 2
y = (5+3)/2 = 4
Center at (2,4)
------
2nd: Find the radius
r = sqrt[(2--3)^2 + (4-5)^2' = sqrt[25+1] = sqrt(26)
------
Equation:
(x-2)^2 + (y-4)^2 = 26
=============================
Cheers,
Stan H.
===============